Please help me solve this geometry question: Determine the equation of the line through the center of the circle that is perpendicular to the diameter with endpoints at (4,2) and (4,-8). It needs to be in slop-intercept form.
Answer by mananth(16949)

End points of diameter = (4,2) and (4,-8).
mid point of these points = centre
Mid point = (4+4)/2 = 4, (-8+2)/2 =-3
Mid point (4,-3)
Slope of diameter = (-8-2)/0
slope of this line is not defined. so it is a line parallel to Y axis
passes through (4,3)
so the equation will be x=4